OpenStudy (anonymous):

Rational Expressions: (2)/(5x) = (-3)/(15)+(1)/(x) These are fractions !

myininaya (myininaya):

\[\frac{2}{5x}*5x=\frac{-1}{5}*5x+\frac{1}{x}*5x\] 2=-x+5 2-5=-x -3=-x x=3

myininaya (myininaya):

-3/15 i reduced it to -1/5 then i multiply both sides by 5x

OpenStudy (anonymous):

Okay, good. We got the same thing ! I just needed it to be solved by someone else. Thanks.